What should be the ratio of the radii of the orbits of electron in Na+10 and hydrogen atom?

Grade:12th pass

1 Answers

Avni Chauhan
5 years ago
radii = 0.529 n2/z
where n = number of shell
and z = atomic number

ratio of Na+10/ H = (1/11) / (1/1)

therefore ratio of radii = 1/11